OpenStudy (anonymous):

Adding and subtracting Rational Expressions 1. 5x-2/ 4x - x-2/4x a. 1 b. x-1/ x c. 0 d. 3/2 5. what is the sum 1/2b + b/2 a. b+1/ 2b+2 b. 2b c. 1/4 d. b^2+1/2b

OpenStudy (anonymous):

1. \[\frac{ 5x-2 }{ 4x } - \frac{ x-2 }{ 4x }\] 5. \[\frac{ 1 }{ 2b } + \frac{ b }{ 2 }\]